The top three members of the management team of an electronic firm purchased the company with their own personal fund and $100 million borrowing. The interest on the loan was 11% and the loan is to be paid off annually by $25 million so that by the end of year 4 the loan will be fully paid off. The unlevered cost of equity for the firm is estimated at 12% and the tax rate is 35%. The free cash flows are estimated for the next four years as follows: $50 million in year 1, $55 million in year 2, $58 million in year 3, and $60 million in year 4. The cash flow is expected to grow at 3% per year after the fourth year.

Estimate the value of the firm as of year 0 using the APV valuation method.
Answer #1
[$ in millions] 0 1 2 3 4 5------->
FCF 50.00 55.00 58.00 60.00
PVIF at 12% 0.89286 0.79719 0.71178 0.63552
PV at 12% 44.64 43.85 41.28 38.13
Sum of PV of FCF - t1 to t4 167.90
Continuing value of FCF = 60*1.03/(0.12-0.03) = 686.67
PV of continuing value = 686.67*0.63552 = 436.39
Unlevered value of FCF 604.30
Interest expense:
Beginning loan balance 100.00 75.00 50.00 25.00
Tax shield on interest at 35% 35.00 26.25 17.5 8.75
PVIF at 11% 0.90090 0.81162 0.73119 0.65873
PV of interest tax shield 31.53 21.31 12.80 5.76
Sum of PV of ITS - t1 to t4 71.40
APV = 604.30+71.40 675.69